OpenStudy (anonymous):

how do i prove this using trig identity: (2sinx+3cosx)^2+(3sinx-2cosx)^2=13

OpenStudy (anonymous):

i did this on my calculator type it in your calulator just as your type on here tell me what you get

OpenStudy (anonymous):

i have to prove it using trigonometric identity without a calculator

OpenStudy (anonymous):

First, use foil to expand them. Then use the following identity \[ \sin^2(\theta)+\cos^2(\theta)=1 \]

OpenStudy (anonymous):

what do i do with the leading coefficients?

OpenStudy (anonymous):

Factor them out.

OpenStudy (anonymous):

What did you get?